dddarryl:

what volume of CO2 will be produced if 2.90 moles of Fe is produced?

justjm:

Is this at STP?

justjm:

We'll need to assume the reaction occurs at STP. \(2.90\ mols\ Fe\cdot\frac{3\ mol\ CO_{2}}{2\ mol\ Fe}\cdot\frac{22.4\ L}{1\ mol}=\) That will be your answer.

dddarryl:

Big help. It is what is over what that confuses me. Thanks a lot!